Calculate the percent ionization of ascorbic acid (HC6H7O6) in solutions of each of the following concentrations (Ka = 8.0e-05.)



(a) 0.185 M

%



(b) 0.575 M

%



(c) 0.740 M

%
Answer #1

a)

HC6H7O6 dissociates as:

HC6H7O6 -----> H+ + C6H7O6-

0.185 0 0

0.185-x x x

Ka = [H+][C6H7O6-]/[HC6H7O6]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((8*10^-5)*0.185) = 3.847*10^-3

since x is comparable c, our assumption is not correct

we need to solve this using Quadratic equation

Ka = x*x/(c-x)

8*10^-5 = x^2/(0.185-x)

1.48*10^-5 - 8*10^-5 *x = x^2

x^2 + 8*10^-5 *x-1.48*10^-5 = 0

This is quadratic equation (ax^2+bx+c=0)

a = 1

b = 8*10^-5

c = -1.48*10^-5

Roots can be found by

x = {-b + sqrt(b^2-4*a*c)}/2a

x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 5.921*10^-5

roots are :

x = 3.807*10^-3 and x = -3.887*10^-3

since x can't be negative, the possible value of x is

x = 3.807*10^-3

% dissociation = (x*100)/c

= 3.807*10^-3*100/0.185

= 2.058 %

Answer: 2.06 %

B)

HC6H7O6 dissociates as:

HC6H7O6 -----> H+ + C6H7O6-

0.575 0 0

0.575-x x x

Ka = [H+][C6H7O6-]/[HC6H7O6]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((8*10^-5)*0.575) = 6.782*10^-3

since x is comparable c, our assumption is not correct

we need to solve this using Quadratic equation

Ka = x*x/(c-x)

8*10^-5 = x^2/(0.575-x)

4.6*10^-5 - 8*10^-5 *x = x^2

x^2 + 8*10^-5 *x-4.6*10^-5 = 0

This is quadratic equation (ax^2+bx+c=0)

a = 1

b = 8*10^-5

c = -4.6*10^-5

Roots can be found by

x = {-b + sqrt(b^2-4*a*c)}/2a

x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 1.84*10^-4

roots are :

x = 6.742*10^-3 and x = -6.822*10^-3

since x can't be negative, the possible value of x is

x = 6.742*10^-3

% dissociation = (x*100)/c

= 6.742*10^-3*100/0.575

= 1.173 %

Answer: 1.17 %

C)

HC6H7O6 dissociates as:

HC6H7O6 -----> H+ + C6H7O6-

0.74 0 0

0.74-x x x

Ka = [H+][C6H7O6-]/[HC6H7O6]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((8*10^-5)*0.74) = 7.694*10^-3

since x is comparable c, our assumption is not correct

we need to solve this using Quadratic equation

Ka = x*x/(c-x)

8*10^-5 = x^2/(0.74-x)

5.92*10^-5 - 8*10^-5 *x = x^2

x^2 + 8*10^-5 *x-5.92*10^-5 = 0

This is quadratic equation (ax^2+bx+c=0)

a = 1

b = 8*10^-5

c = -5.92*10^-5

Roots can be found by

x = {-b + sqrt(b^2-4*a*c)}/2a

x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 2.368*10^-4

roots are :

x = 7.654*10^-3 and x = -7.734*10^-3

since x can't be negative, the possible value of x is

x = 7.654*10^-3

% dissociation = (x*100)/c

= 7.654*10^-3*100/0.74

= 1.034 %

Answer: 1.03 %